OpenStudy (anonymous):

Solve: 2(4x-3) – 8 = 4 + 2x

OpenStudy (anonymous):

Do you know what the distributive property is?

OpenStudy (anonymous):

yes it's going to be 8x-6-8=4+2x

OpenStudy (anonymous):

Right. So now you're gonna want to get all of the x terms on one side and all the normal terms on the other.

OpenStudy (anonymous):

Combine like terms on the left side. Add 14 to both sides. Subtract 2x from both sides.

OpenStudy (anonymous):

Right, good job. So what will it look like now?

OpenStudy (anonymous):

8x-2x=4+14

OpenStudy (anonymous):

Right, which is just \(6x=18\) Now all you have to do is divide to get x alone.

OpenStudy (anonymous):

10/6 is reduced to 3

OpenStudy (anonymous):

Correct! So \(x=3\) :)

OpenStudy (anonymous):

yes thanks can yhu help with another one

OpenStudy (anonymous):

Solve: 5(1 - 5x) + 5(-8x + 2) = -4x – 8x. Round your answer to four decimal places.